### v1

# EVINCE: Optimizing Adversarial LLM Dialogues via Conditional Statistics and Information Theory

Abstract
This paper introduces EVINCE (Entropy and Variation IN Conditional Exchanges), a dialogue framework advancing Artificial General Intelligence (AGI) by enhancing versatility, adaptivity, and reasoning in large language models (LLMs). Leveraging adversarial debate and a novel dual entropy theory, EVINCE improves prediction accuracy, robustness, and stability in LLMs by integrating statistical modeling, information theory, and machine learning to balance diverse perspective exploration with strong prior exploitation. The framework’s effectiveness is demonstrated through consistent convergence of information-theoretic metrics, particularly improved mutual information, fostering productive LLM collaboration. We apply EVINCE to healthcare, showing improved disease diagnosis, and discuss its broader implications for decision-making across domains. This work provides theoretical foundations and empirical validation for EVINCE, paving the way for advancements in LLM collaboration and AGI development.
EVINCE: Optimizing Adversarial LLM Dialogues
via Conditional Statistics and Information Theory
Edward Y. Chang Computer Science Clinical Mind AI Stanford University echang@cs.stanford.edu Draft: 2/15/2024; Revised: 8/15/2024
1 Introduction
The pursuit of Artificial General Intelligence (AGI) remains a central goal of AI research. We propose a paradigm shift in this quest: utilizing multiple Large Language Models (LLMs) engaged in synergistic dialogues as a crucial step towards AGI. This approach, we contend, addresses key limitations of current AI systems and provides a novel pathway to more robust, versatile, and capable artificial intelligence. Specifically, our work targets three core AGI characteristics: versatility, adaptivity, and reasoning capability.
Current LLMs, despite their remarkable capabilities, face significant challenges, including hallucination (generating false or nonsensical information), bias (reflecting and potentially amplifying societal prejudices), and limited reasoning (difficulties in complex problem-solving and logical inference). We posit that multi-agent dialogue systems offer a promising avenue to address these challenges. By fostering diversity and debate among LLMs, these systems can mitigate biases and promote enhanced reasoning capabilities. Furthermore, the iterative nature of multi-round dialogues allows for continuous context enrichment, enabling LLMs to access more precise information and formulate more accurate responses, thus reducing the occurrence of hallucinations.
Our previous work, SocraSynth Chang (2023a), addresses LLM limitations through structured multi-agent dialogues. By leveraging both adversarial and collaborative interactions between LLMs, SocraSynth demonstrates quantifiable improvements across various domains. In healthcare, it achieves a 5% improvement in diagnostic accuracy Chang and et al. (2023), while in news analysis, it identifies and explains biases in articles Chang (2024c). These advancements showcase enhanced versatility, adaptivity, and reasoning capability—key AGI characteristics.
SocraSynth’s adversarial component promotes the exploration of diverse perspectives, while its collaborative component fosters rigorous reasoning to reach well-reasoned conclusions. This synergy has yielded measurable gains beyond healthcare and bias mitigation, extending to geopolitical analysis Chang (2023b), corporate planning Tsao (2023), investment banking Chang (2024b), and emotional behavior modeling Chang (2024a). These results demonstrate SocraSynth’s effectiveness in mitigating LLM limitations and achieving substantial performance improvements across various applications, highlighting its potential for advancing towards AGI’s generalized problem-solving capabilities.
Although SocraSynth has demonstrated effectiveness, it relies on “contentiousness” as a qualitative measure to moderate the linguistic behaviors of participating LLMs. For instance, a high contentiousness value (0.9 out of 1.0) might lead LLMs to challenge each other’s assumptions and propose alternative solutions, while a low value (below 0.3) could encourage them to synthesize their viewpoints and find common ground. While the concept of contentiousness has proven useful in guiding SocraSynth dialogues, its qualitative nature limits its precision and explainability.
In this paper, we extend SocraSynth by introducing three theoretical pillars to quantify “contentiousness” and moderate dialogues based on statistical and information theories. These pillars, collectively referred to as EVINCE (Entropy and Variation IN Conditional Exchanges), provide quantitative measures for justifiable and explainable multi-agent dialogue moderation and evaluation.
These pillars are:
  1. 1.
    Inclusiveness Exploration: We develop methods to ensure dialogues explore all potential perspectives. We use conditional statistics via in-context learning Xie et al. (2021) to “free” an LLM agent from its default “maximal likelihood” next-token prediction behavior, allowing it to adopt specific stances. We introduce a dual entropy optimality theory to balance the exploration of new ideas with adherence to priors, thus optimizing information exchange between agents for comprehensive and stable discourse.
  2. 2.
    Information Flow Dynamics: We introduce information theory-based metrics to quantify and optimize dialogue dynamics. These measure information diversity (entropy), novelty (statistical divergence scores), and inter-agent persuasion (mutual information). These metrics enable us to assess and enhance the quality and efficiency of information flow within the multi-agent system, fostering rich and productive exchanges.
  3. 3.
    Reasoning Quality and Coherence: We establish frameworks to assess the logical structure and coherence of multi-agent reasoning. This pillar evaluates argument validity, analytical depth, and dialogue coherence. We synergistically integrate our previously developed CRIT system Chang (2023c) into SocraSynth. CRIT, which combines Socratic methods with formal reasoning techniques, enhances our ability to conduct critical thinking through evaluating argument quality, information-source credibility, and overall “reasonableness” within the dialogue. This integration ensures that the collective reasoning of LLM agents is not only diverse but also logically sound and aligned with the dialogue’s objectives.
The core strength of SocraSynth and EVINCE in advancing towards AGI lies in their ability to enhance key AGI characteristics through multi-agent dialogues. By employing conditional statistics and information theory, they boost versatility and adaptivity, allowing LLMs to transcend their typical “maximal likelihood” behaviors and mimic how humans adapt their linguistic behaviors to complete tasks. The framework’s debate structure fosters advanced reasoning and critical thinking capabilities. Together, these advancements represent significant steps towards achieving the complex, intricate capabilities required for AGI.
The main contributions of this paper are:
  1. 1.
    EVINCE Framework Design: Unlike approaches that use debate merely to improve accuracy via redundancy, EVINCE facilitates information discovery, bias mitigation, and decision-making that requires both breadth and depth of information.
  2. 2.
    Theoretical Foundations: We establish a theoretical basis for EVINCE, rooted in conditional Bayesian statistics, mutual information, and dual entropy. These principles are applied to measure, monitor, and modulate collaborative LLM interactions, contributing to a deeper understanding of how LLMs can effectively cooperate for improved decision-making. The dual entropy theory is novel and ground-breaking, illustrating how a productive decision-making process should start with room for diverse input and stable objectives, and then, through information exchange, converge to optimal decision/prediction.
  3. 3.
    Empirical Validation: We provide empirical validation of EVINCE’s underlying theories, highlighting the framework’s effectiveness in balancing exploration and exploitation to enhance prediction accuracy. We also introduce a set of maxims derived from our empirical findings, offering practical guidance for optimizing mutual information and minimizing various divergence measures.
The remainder of this paper is structured as follows: Section 2 discusses related work and addresses common misconceptions about our approach. Section 3 presents the three pillars of EVINCE, including its maxims, theories, and algorithms. Section 4 provides empirical evidence of EVINCE’s benefits by revisiting and expanding upon previous SocraSynth experiments in healthcare and bias detection. Section 5 offers concluding remarks and discusses future directions.
2 Related Work
The core objective of adversarial debate, as embodied in SocraSynth with EVINCE, is to foster diverse opinions and challenge assumptions, ultimately leading to more comprehensive and informed decision-making. This contrasts with traditional ensemble learning methods, which prioritize error diversity for improved accuracy. This section surveys ensemble learning and adversarial learning, highlighting the unique aspects of EVINCE and its theoretical underpinnings.
2.1 Ensemble Learning
Ensemble methods, such as Bagging Breiman (1996), Boosting Freund and Schapire (1997), and Mixtures of Experts Jacobs et al. (1991), traditionally aim to enhance machine learning performance by combining predictions from multiple models to achieve error diversity. Similarly, early LLM debate frameworks, e.g., Michael et al. (2023); Chan et al. (2023); Liang et al. (2023); Du et al. (2023) have focused primarily on improving answer accuracy rather than fostering a comprehensive exploration of information for critical decision-making.
SocraSynth distinguishes itself from traditional ensemble methods by prioritizing the generation of diverse predictions over the mere avoidance of errors. This is achieved through a dynamic protocol that adaptively adjusts the “contentiousness” level of the debate, encouraging models to initially explore a wide range of perspectives and rigorously assess the quality of arguments. Through constructive exchange of ideas in subsequent iterations, SocraSynth converges on a more informed and robust decision. Its approach, grounded in EVINCE’s dual entropy theory, allows for the exploration of various ideas by one LLM while maintaining stability through the other, effectively preventing premature convergence on a narrow set of perspectives.
2.2 Addressing Common Misconceptions about Adversarial LLM Dialogue
The rapid growth in academic conference submissions has led to an expanded reviewer pool, potentially impacting review quality. In this context, some misconceptions about our work have arisen, which we address here:
  1. 1.
    Misconception: Multiple LLMs cannot mitigate hallucination.
    Some reviewers have expressed skepticism about the ability of multiple LLMs to reduce hallucination. However, our approach does not simply aggregate potentially flawed outputs. Instead, SocraSynth leverages the strengths of multiple models to cross-check, challenge, and refine information. This process is analogous to peer review in scientific discourse, where collective scrutiny enhances the reliability of conclusions.
  2. 2.
    Misconception: Multiple LLM dialogues cannot lead to knowledge discovery.
    Despite evidence provided in Chang (2023a) and Chang and Chang (2023), some reviewers remain skeptical about the potential for knowledge discovery through LLM dialogues. Our work demonstrates that carefully structured interactions between LLMs can indeed lead to novel insights and knowledge synthesis by leveraging the polydisciplinary knowledge representation inherent in LLMs.
  3. 3.
    Misconception: The approach lacks theoretical foundation.
    Our approach is grounded in conditional statistics and information theory, which we use to optimize LLM interactions. This theoretical foundation ensures that the dialogue converges towards more accurate and less biased outcomes, addressing concerns about the rigor of our methodology.
By addressing these misconceptions, we aim to clarify the scientific basis and potential of our approach. The synergistic interaction between LLMs in SocraSynth, coupled with EVINCE’s theoretical pillars, represents a novel method for enhancing the capabilities of language models. This approach has applications in reducing bias, mitigating hallucination, and potentially advancing the field towards more robust AI systems.
Metric Pros Cons Remedies
Cross Entropy (CE) Shore and Johnson (1980) Measures how well the predictions of one model fit the actual distribution of another model’s outputs (asymmetric). Computationally intensive especially with large models and data sets; sensitive to the exact nature of probability distributions. Optimize computation strategies; use approximations or sampling methods to manage large data sets or complex models.
Entropy Shannon
Shannon (1948)
Indicates level of diversity; high suggests exploration of possibilities, and low for confidence on few choices High entropy might indicate noise rather than useful diversity; low entropy might mask important variability. Use critical reading methods (Appendix A) to assess argument quality; implement noise detection to differentiate between useful diversity and noise.
Jensen-Shannon
Divergence (JS)
Lin (1991)
Symmetric and bounded (0 to 1), providing an interpretable measure of distributional differences. May be less sensitive to small differences between distributions. Increase sensitivity settings or resolution of the metric; combine with other metrics to capture finer distinctions between distributions.
KL Divergence
Kullback (1951)
Measures difference between two probabilistic distributions. Asymmetric; not well-defined if a distribution has zero probabilities Use smoothing techniques to avoid zero probabilities; consider symmetric alternatives like JS divergence
Mutual Info (MI) Shore and Johnson (1980) Measures reduction of uncertainty; symmetric. Does not indicate the directionality of information flow. Supplement with directional information metrics; normalized with max entropy of A and B.
Wasserstein
Distance (WD)
Kantorovich (1942)
Direct measure of how similar or different the model outputs are; it depicts symmetric relationship. Not bounded but can be normalized or bounded for consistent interpretation. Define context-specific bounds for low, medium, and high divergence; consider normalizing it for non-directional comparisons.
Table 1: Summary of metrics for assessing LLM debates (equations are presented in Appendix D)
3 Algorithm, Maxims, and Theories
Problem Statement:
Organize a structured debate between two equally competent large language models (LLMs), LLMA and LLMB, to conduct $t$ rounds. At each round $t$, each model produces a probability distribution, denoted as $P_{A}^{(t)}$ and $P_{B}^{(t)}$, over $C$ possible outcomes, accompanied by supporting arguments $R_{A}^{(t)}$ and $R_{B}^{(t)}$. The goal is to design an iterative debate process that leverages the structured exchange of arguments to enable the models to converge on an optimal prediction distribution $P^{*}$ across the $C$ classes. Optimality is defined as achieving the highest possible accuracy on the prediction task.
To address this problem, we first introduce key metrics from information theory that form the basis of our approach.
3.1 Preliminaries in Information Theory
This section presents the key metrics used to measure information diversity, similarity, divergence, and other relevant factors within EVINCE. These metrics serve three primary objectives:
3.1.1 Fostering diversity of perspectives while ensuring reasoning quality
  • Wasserstein Distance (WD): Measures distribution difference between predictions to identify exploration opportunities Kantorovich (1942); Rubner et al. (2000); Villani (2008).
  • Shannon Entropy or Relative Entropy: Measures diversity of perspectives Cover and Thomas (2006); Shannon (1948).
  • Reasoning Quality: The CRIT algorithm evaluates the logical soundness and persuasiveness of supporting arguments Chang (2023c), helping to identify and mitigate hallucinations and poorly-reasoned arguments.
3.1.2 Exploring new possibilities while adhering to the dialogue subject
  • Correlation Coefficients: Tracks the evolution of opinions and assesses debate stability Brown et al. (2005) toward the goal of the dialogue.
  • Mutual Information (MI): Quantifies information overlap to ensure focused and productive debates Cover and Thomas (2006) and measures the degree of agreement/disagreement.
3.1.3 Examining information convergence and establishing termination criteria
  • Jensen-Shannon (JS) Divergence: Assesses similarity between probability distributions Lin (1991) (symmetric).
  • Cross Entropy (CE): Measures the asymmetric difference between prediction distributions Shore and Johnson (1980).
  • Kullback-Leibler (KL) Divergence: Reveals asymmetric differences between probability distributions Kullback (1951).
Table 1 summarizes these metrics and their pros and cons. Next, we present the EVINCE algorithm and explain how these metrics are used to moderate an LLM dialogue to achieve a balance between exploration and exploitation, leading to optimal prediction outcomes. In Section 3.3, we will introduce several maxims and prove the dual entropy theorem, which provides a theoretical foundation for the EVINCE algorithm.
[…]
Figure 1: Specifications of Algorithm EVINCE. Key points: 1) Asymmetric Start: In Step #1, LLMA initiates with opening arguments based solely on the given information, while LLMB starts with access to LLMA’s prediction and arguments for refutation. 2) Termination Criteria: The while loop in Step #2 considers three factors: Wasserstein distance, mutual information, and argument quality. EVINCE terminates if the dialogue ceases to make significant progress. 3) Further Details: Maxims #1 to #4 provide additional explanations of the algorithm’s principles. 4) Argument Evaluation: For details on the CRIT algorithm used to assess argument quality, refer to Chang (2023c).
3.2 EVINCE Algorithm Specifications
Figure 1 formally specifies the detailed operations of EVINCE.
Setup and Initialization:
EVINCE employs two equally competent LLM instances, $LLM_{A}$ and $LLM_{B}$, which can be different models (e.g., GPT and Claude) or independent instances of the same model. Given an information set $S$ and a class-label set $C$, EVINCE outputs a probability distribution over $C$ with justifications. For example, $S$ could represent a patient’s symptoms and $C$ a set of diseases, with EVINCE moderating a dialogue to predict the patient’s disease(s).
The debate “contentiousness” is initially set to 90% to encourage disagreement and explore diverse perspectives. This high level is maintained for the first two iterations, then gradually reduced by $x\%$ in each subsequent iteration to shift from exploration to exploitation [^1] The setting of $x$ at the beginning of a dialogue iteration can depend on the values of all moderation subroutines. Our empirical experience suggests $x=25\%$ because a typical dialogue concludes in five iterations, and $x=20\%$ would set contentiousness to be 90%, 90%, 68%, 50%, and 38%.. Table 2 demonstrates how GPT adapts its linguistic behavior based on the “contentiousness” parameter, showcasing the model’s ability to adjust to different contextual settings. Importantly, this table is generated by GPT itself, not by human moderators, as detailed in Chang (2023a) and further explored in Section 4.
C.L. Tone Emphasis Language
$0.9$ Highly confrontational; focused on raising strong ethical, scientific, and social objections. Flagging risks and downsides; ethical quandaries, unintended consequences, and exacerbation of inequalities. Definitive and polarizing, e.g., “should NOT be allowed,” “unacceptable risks,” “inevitable disparities.”
$0.7$ Still confrontational but more open to potential benefits, albeit overshadowed by negatives. Acknowledging that some frameworks could make it safer or more equitable, while cautioning against its use. Less polarizing; “serious concerns remain,” “needs more scrutiny.”
0.5 Balanced; neither advocating strongly for nor against gene editing. Equal weight on pros and cons; looking for a middle ground. Neutral; “should be carefully considered,” “both benefits and risks.”
0.3 More agreeable than confrontational, but maintaining reservations. Supportive but cautious; focus on ensuring ethical and equitable use. Positive but careful; “transformative potential,” “impetus to ensure.”
0.0 Completely agreeable and supportive. Fully focused on immense potential benefits; advocating for proactive adoption. Very positive; “groundbreaking advance,” “new era of possibilities.”
Table 2: Changes in linguistic behaviors of LLMs at different contentiousness levels.
Moderation Subroutines:
EVINCE employs four key subroutines to manage the exploration-exploitation tradeoff, ensuring information diversity, quality, and stability:
  • CRIT: Evaluates argument quality and source credibility. Low scores can trigger dialogue termination.
  • WD (Wasserstein distance): Assesses prediction diversity, expected to decrease over time.
  • MI (Mutual Information): Evaluates dialogue convergence. Stagnation can trigger termination.
  • Additional metrics: KL divergence, Jensen-Shannon divergence, and cross entropy ensure evaluation consistency (see Table 1).
When all metrics reach a plateau, indicating no further improvement, EVINCE terminates the dialogue. At this point, both LLM instances are prompted to jointly deliver a conciliatory conclusion with comprehensive arguments and counterarguments. Additionally, they provide a list of missing information that, if obtained, could potentially improve prediction accuracy and reliability.
Dialogue Iterations (Steps 1 and 2):
Given the contentiousness level set for a dialogue iteration, each LLM generates a new prediction distribution on $C$ with arguments, based on the other’s last prediction and accumulated reasons as the context.
Final Output (Step 3):
  • Combine final predictions from both LLMs.
  • Weight predictions based on supporting argument quality (evaluated by CRIT).
  • Handle uncertainty by soliciting missing information when final prediction entropy is high (see Chang (2024b) for RAG application).
EVINCE creates a structured debate between AI models, balancing diverse viewpoints with consensus-building to produce accurate, well-reasoned predictions.
3.3 Maxims with Theoretical Foundations
Progress towards the optimality goal is guided and measured by metrics introduced in Section 3.1. This section explains how these metrics can be used in complementary ways to facilitate proper trade-offs between diversity and convergence, exploration and exploitation, and several other factors. In the EVINCE algorithm presented in Figure 1, we have annotated the steps to which these four maxims are applied.
Maxim #1: Orchestrate Two Equally Competent LLMs in Structured Debate:
Integrating two equally competent LLMs ensures a balanced exchange of insights and avoids bias. This adversarial setup fosters diversity in predictions, each supported by justifications, promoting critical evaluation and uncovering potential blind spots.
How? Choosing LLMs with comparable performance on a shared validation set, a balanced debate can be ensured. Suitable models include GPT-4, Claude, and Gemini. Conditioning different instances of the same LLM to support opposing stances on a subject matter can also be effective due to the theoretical justification of in-context learning with conditional Bayesian statistics Xie et al. (2021).
Maxim #2: Encourage the Accurate Rather Than the “Popular” Prediction:
Typically, LLMs, with their maximum likelihood next-token prediction objective, tend to favor the most popular predictions. By conditioning LLMs within specific contexts, we can prioritize specific stance over popularity, mitigating confirmation biases.
How? Using the proxy metrics in Table 1, EVINCE dynamically adjusts the “contentiousness” level in debates. These metrics quantify agreement, diversity, and mutual information, promoting productive information exchange and enhancing prediction quality.
Maxim #3. Evaluating the Convergence Rate of the Predictions Across the Iterations:
This aspect focuses on measuring how quickly and effectively the predictions from the LLMs converge over successive rounds, assessing the efficiency of the debate and aggregation mechanisms.
How? Convergence is assessed by measuring mutual information and using proxy metrics such as Wasserstein distance and cross entropy. When the mutual information is high or the similarity between predictions is high, the debate is considered to be converging.
Maxim #4. Combine Predictions Weighted by Diversity and Quality:
Weighting the probability distributions from two LLMs based on diverse probabilistic insights and argument quality.
How? Following these three maxims:
  • Maxim #4.1 Prediction Reliability: Estimate the reliability of predictions using entropy-based measures to quantify uncertainty and information content. Typically, lower entropy indicates higher confidence in a prediction, suggesting higher reliability.
  • Maxim #4.2 Argument Quality: Evaluate the quality of supporting arguments using techniques inspired by the Socratic method. This includes identifying logical fallacies, assessing the relevance and credibility of evidence.
  • Maxim #4.3 Aggregation: Employ a weighted aggregation method, such as a Bayesian model to combine weighted predictions accounting for both probabilistic insights and the quality of supporting arguments.
3.4 Entropy Duality Theorem (EDT)
Theorem EDT: Optimal Pairing of LLMs for Probabilistic Prediction Accuracy.
The optimal pairing of LLMs for diagnosis accuracy, in terms of stability, accuracy, and robustness, occurs when the LLMs are 1) equivalent in the quality of the information they process, and 2) exhibit contrasting entropy values in their prediction distributions—one high and one low.
Proof.
Please see Table 3. ∎
Theorem EDT: Optimal Pairing of LLMs for Probabilistic Prediction Accuracy.
The optimal pairing of LLMs for diagnosis accuracy, in terms of stability, accuracy, and robustness, occurs when the LLMs are 1) equivalent in the quality of the information they process, and 2) exhibit contrasting entropy values in their prediction distributions—one high and one low.
[Proof]:
[…]
Step 6: Conclude the proof. By combining the prediction distributions of LLMA and LLMB, with one having high information entropy and the other low, we achieve an optimal balance that maximizes the lower bound of the combined information entropy. This balance between exploration (high entropy) and exploitation (low entropy) optimizes overall prediction accuracy and robustness, as proved through information theory and the properties of entropy. Thus, the theorem is established. $\square$
Table 3: Proof of Entropy Dual Theorem (EDT)
4 Empirical Study
This empirical study investigates the application of EVINCE to disease diagnosis, leveraging large language models (LLMs) as diagnostic tools. We aim to validate the following three hypotheses:
  1. 1.
    Contentiousness & Prediction Quality: Initial LLM disagreement (measured by Wasserstein distance) increases with higher initial contentiousness but decreases as debate progresses. Individual LLM prediction uncertainty (Shannon entropy) will follow a similar pattern.
  2. 2.
    EDT Effectiveness & Confusion Matrices: LLM pairs following the Entropy Duality Theorem (EDT) will have complementary error patterns, leading to higher combined prediction accuracy than non-EDT pairs.
  3. 3.
    EVINCE & Historical Misdiagnoses: EVINCE, applied to real-world data, will improve diagnostic accuracy and identify potential misdiagnoses or ambiguities within the ground truth.
(a) GPT4 pairs Claude
(b) GPT4 pairs Gemini
Figure 2: Pre-/post-debate accuracy on all patients on all diseases shows EVINCE helps
Problem Statement:
Given a set of symptoms $S$ and a context $\kappa$, the goal is to predict a probability distribution of top-$k$ diseases over a set of $C$ possible diseases. This is represented as $P=\text{LLM}(S,\kappa)$, where each LLM generates top-$k$ predictions on $C$ ($k\leq|C|$) based on the input symptoms $S$ and context $\kappa$.
$P=\left(p(\text{top}\,1\,\text{to}\,k\in C\mid S,\kappa)\right).$
Context $\kappa$ is where dual entropy is adjusted through three parameters: temperature, the value of $k$ in top-$k$, and the contentious level $\Delta$. A distribution tends to have high entropy when all three parameters are set high, and vice versa.
Resources, Dataset & Data Preparation:
Our study utilizes a dataset from Kaggle Patil (2020), comprising 4,921 patient records. Each record includes the diagnosed disease and up to 17 symptoms (e.g., fever, cough, fatigue, itchiness, difficulty breathing). After removing duplicates, we obtain 304 unique diagnostic instances spanning 40 diseases. (The refined dataset is provided as supplementary data.) Each instance serves as a test case where EVINCE leverages the inherent knowledge of LLMs (GPT-4, Gemini, and Claude3) without training them through few-shot techniques on this specific dataset. Our computing resources are sponsored by Azure through a Stanford grant.
Evaluation:
We assess prediction quality using the top-k Mean Reciprocal Rank (MRR). If one of the top-k predicted diseases matches the ground truth diagnosis, the score is the reciprocal of its rank (1 for the top prediction, 1/2 for the second, 1/3 for the third, etc.). If none of the top-k predictions are correct, the score is 0.
4.1 Study #1: Post vs. Pre-Debate Accuracy
We employed GPT-4, Gemini, and Claude3 to perform independent disease predictions on 304 patient instances, then used EVINCE to pair them and evaluate performance gain.
Experimental Setup:
We set $k=5$ for both LLM agents, with one agent at high temperature and the other at low temperature. The contentiousness level was set very high ($\Delta=0.9$ out of 1) to encourage significant cross entropy. Setting $k=5$ ensures some minimal common ground, fostering meaningful interaction. High contentiousness promotes counterarguments and information exchange.
Pre- and Post-Debate Evaluation:
We conducted two sets of experiments:
1. Constrained Prediction (Baseline): We limited disease predictions to the 40 labels in the dataset, mimicking common supervised learning assumptions. This yielded high accuracy (95-97%) but is unrealistic for real-world diagnosis where general practitioners consider all possibilities. This constraint highlights LLMs’ flexibility, as they’re less prone to overfitting erroneous labels (further discussed in subsequent studies).
2. Unconstrained Prediction: We removed the label constraint to better simulate real-world conditions. All 304 patient cases yielded stable results across GPT-4, Gemini-3, and Claude-3, with a standard deviation of just 1.5%. Prior to debate (light blue bars in Figure 2), GPT-4 led in accuracy (82.8%), followed by Gemini (80.3%) and Claude (79.5%).
EVINCE Performance:
Implementing EVINCE with GPT-4 and Claude-3 pairing and GPT-4 and Gemini-3 pairing consistently improved accuracy by 4-5 percentage points (green bars in Figure 2). The GPT-4 and Claude-3 pairing achieved 87.5% accuracy (Figure 2(a)), rivaling state-of-the-art clinical performance like the REFUEL algorithm Peng et al. (2018).
Discussion:
The remaining 12.5% inaccuracy for the GPT-Claude pairing might not be solely attributable to EVINCE. Considering the potential 11% US misdiagnosis rate reported by Johns Hopkins Newman-Toker et al. (2023b), this discrepancy could indicate mislabeled data in the original dataset. This presents a groundbreaking opportunity: EVINCE could potentially identify and correct errors in existing datasets, a concept we explore further in Section 4.3.
4.2 Study #2: Confusion vs. Opportunities
(a) GPT liver c-matrix
(b) Claude liver c-matrix
Figure 3: Confusion matrices
(a) Entropy
(b) WD %
(c) Norm. MI
Figure 4: Entropy, WD, and normalized MI
Two primary factors contribute to EVINCE’s improved diagnostic accuracy: (1) structured debates with reasoning encourage LLMs to explore alternative diagnoses in both breadth and depth, leading to more comprehensive analysis and decision-making (see Appendices B and C); and (2) pairing high- and low-entropy LLMs balances exploratory diversity with exploitative stability, resulting in more robust and high-quality decisions, as demonstrated in this second study.
Analysis of Confusion Matrices:
We use confusion matrices to analyze two LLMs’ performance in diagnosing Hepatitis types A to E (Fig. 3):
  • GPT-4 shows limited accuracy, particularly for types C and D, achieving only 50% accuracy for types A and B.
  • Claude exhibits a wider spread of predictions across all Hepatitis types.
These matrices highlight how Claude’s flexibility in exploring diverse diagnostic hypotheses aids the debate process. Claude’s initial uncertainty (high entropy) brings new information, potentially challenging and correcting GPT-4’s more confident (low entropy) predictions. This dynamic interplay exemplifies EVINCE’s balance between exploration and exploitation, leading to more accurate and comprehensive diagnoses.
Observations from Information Metrics:
  • Entropy Stabilization: Figure 4(a) shows entropy levels of both LLMs stabilizing after three debate rounds, indicating convergence towards a similar, stable entropy state.
  • Wasserstein Distance Improvement: Figure 4(b) demonstrates consistent improvement in Wasserstein distance (WD) between the two models’ predictions over successive rounds.
  • Mutual Information Increase: Figure 4(c) reveals a 14% improvement in normalized mutual information (MI) between GPT-4 and Claude’s prediction distributions, suggesting increased shared information throughout the debate.
  • Convergence of Divergence Metrics: Figure 5 shows consistent convergence of all divergence metrics.
Comparative Performance:
EVINCE demonstrates a 5% higher accuracy rate in diagnosing specific types of liver diseases compared to a baseline approach (Figure 2(a)), underscoring its capability to handle complex diagnostic scenarios effectively.
Figure 5: Convergence of all metrics
4.3 Study #3: Ground-Truth Remediation
This study illustrates how EVINCE can identify potential misdiagnoses, explain the reasoning behind them, and recommend corrective actions. Traditionally, machine learning scientists rely on labeled data as “ground truth.” However, as evidenced by research like that of Newman-Toker et al. (2021) Newman-Toker et al. (2023a) from Johns Hopkins, misdiagnosis is a widespread issue in healthcare systems globally. These erroneous diagnoses, often treated as ground truth, can be perpetuated by supervised learning algorithms, exacerbating the problem within the healthcare system.
(a) GPT & Claude top-5
(b) WD %
(c) Mutual Info.
Figure 6: Remediation: Jaundice to Hepatitis
In the debate scenario detailed in Appendix C, where Jaundice is the ground truth diagnosis, Figure 6(a) illustrates initial differences between GPT-4 and Claude’s predictions. Jaundice is absent in GPT-4’s top-5 (with 0% in red), while ranked third by Claude. Although Claude influences GPT-4 to include Jaundice in its third prediction in the second round, subsequent rounds see both LLMs drop Jaundice to the fourth position of 10%.
Meanwhile, Hepatitis A, initially GPT-4’s top prediction (30% in dark blue), is quickly demoted to fifth and eventually drops out of the top-5 entirely due to Claude’s influence. Hepatitis B, initially ranked second by GPT-4 and top by Claude, stabilizes in the second position in rounds 3 and 4 (in light blue). Notably, Hepatitis C rises from second place on both lists to the top position and remains there (in black).
Figure 7: Convergence of all metrics
As demonstrated in the previous case study, Wasserstein distance (WD) effectively measures the divergence between LLM predictions and assesses debate convergence. Figures 6(b) and 6(c) show that WD stabilizes after three debate rounds, coinciding with a plateau in normalized mutual information (MI) between GPT-4 and Claude. This stabilization suggests their predictions converge.
Figure 7 illustrates the convergence of all divergence metrics—including Jensen-Shannon divergence, cross-entropy, and Kullback-Leibler divergence—particularly between the second and third rounds. Although the final joint prediction for Hepatitis C reached a high consensus of 37.5%, it deviates from the actual condition of Jaundice, which the Kaggle dataset reports with 10% confidence. EVINCE provides general practitioners with alerts and suggests remedial actions (see Appendices D.9 and C.8) to address this discrepancy. Recommended actions include querying additional symptoms from the patient and conducting specific laboratory tests.
4.4 Experiment Remarks
EVINCE initiates debates with high contentiousness, encouraging dual prediction entropy between LLMs, as supported by the EDT theorem. It utilizes normalized mutual information (MI) to track shared knowledge accumulation throughout the debate, while Wasserstein distance (WD) and Jensen-Shannon divergence (JSD) quantify dissimilarity between LLM predictions.
These metrics (EDT, WD, JSD, MI) provide a comprehensive view of debate progress. WD and JSD assess the potential for further communication and refinement, while MI monitors shared understanding, aiding in determining the optimal stopping point.
The asymmetric nature of KL divergence and cross entropy warrants further investigation. Despite eventual convergence in our case studies, discrepancies observed in the second round (one direction increasing while the other decreases) suggest potential value in exploring asymmetric information. Future work will re-evaluate the use of these metrics if asymmetry proves beneficial.
Besides generating final joint disease predictions, EVINCE provides:
  • Recommendations for additional symptom inquiries and lab tests to improve accuracy.
  • Suggestions to query symptom onset, duration, severity, trends, and associated symptoms (documented in Appendices B.8 and C.9).
These recommendations have been verified by general practitioners to be valuable.
5 Concluding Remarks
This paper introduces EVINCE, a framework designed to facilitate collaborative dynamics among Large Language Models (LLMs) through structured, adversarial debates. Our research demonstrates that EVINCE significantly advances the pursuit of Artificial General Intelligence (AGI) by enhancing three core characteristics: versatility, adaptivity, and reasoning capability. By addressing key limitations of current AI systems, including hallucination, bias, and limited reasoning, EVINCE provides a novel pathway towards more robust and capable artificial intelligence.
The core strength of EVINCE lies in its three theoretical pillars: Inclusiveness Exploration, Information Flow Dynamics, and Reasoning Quality and Coherence. These pillars, grounded in conditional statistics and information theory, enable LLMs to transcend their typical “maximal likelihood” behaviors, mirroring human adaptability in linguistic tasks. The integration of the CRIT system, which combines Socratic methods with formal reasoning techniques, further enhances critical thinking and ensures logically sound and objective-aligned collective reasoning.
Our empirical validation demonstrates EVINCE’s effectiveness in improving prediction accuracy across various domains, notably achieving a 5% improvement in medical diagnosis tasks. The framework has also shown promise in identifying biases in news articles Chang (2024c), showcasing its potential for broader applications in fields such as geopolitical analysis Chang (2023b) corporate planning Tsao (2023), and emotional behavior modeling Chang (2024a).
While EVINCE has demonstrated significant potential, our future work will focus on refining key aspects of the system. We aim to enhance the contentiousness parameter by tying it more rigorously to metrics such as Wasserstein distance, mutual information, and CRIT scores. Additionally, we plan to explore alternative methods for inducing dual-entropy conditions, moving beyond current parameters like temperature and top-k selection.
Comprehensive ablation studies will be crucial in determining optimal configurations for various applications, ensuring EVINCE’s adaptability to diverse contexts and challenges. Through these continued refinements and explorations, we aim to push the boundaries of collaborative AI frameworks, bringing us closer to realizing the full potential of artificial general intelligence.
In conclusion, EVINCE represents a significant step forward in AI research, offering a structured approach to leveraging the collective intelligence of multiple LLMs. By addressing current limitations and enhancing key AGI characteristics, EVINCE paves the way for more versatile, adaptive, and capable AI systems, bringing us closer to the goal of artificial general intelligence.
Acknowledgements
This work has been significantly shaped by insights from several key figures in the field. Microsoft CSO Eric Horvitz’s concept of “polydisciplinary” representation in LLMs was a major influence. Demis Hassabis of DeepMind provided valuable input on end-to-end architecture without human interference. Prof. Stuart Russell’s assertion that LLMs mimic human linguistic behavior to achieve goals was pivotal in our approach to modeling linguistic behaviors and utilizing Bayesian frameworks.
My colleague Prof. Vaughan Pratt’s feedback on the initial paper inspired the use of “contentiousness” to foster adversarial debate, a key component of our framework. I am grateful to Prof. Monica Lam for her collaborative support throughout this research. Additionally, I thank my colleagues at Stanford Clinical Mind AI for their invaluable domain knowledge in disease diagnosis.
Appendix A: Evaluative Phase of EVINCE
EVINCE uses the Socratic method to evaluate the “reasonableness” of a set of arguments that support a subject matter. The Socratic method is a questioning technique used in teaching and philosophy to encourage critical thinking and self-discovery Wikipedia (2023). The method involves asking a series of questions to explore complex ideas and help individuals arrive at their own understanding of a concept. It is based on the belief that knowledge cannot be simply imparted, but must be discovered through a process of questioning and dialogue.
To illustrate how these methods can practically be applied, let’s use the example of critical reading. Critical reading is a crucial component of critical thinking, which involves evaluating the quality and credibility of written materials, from research papers to blog posts Lai et al. (2017); Paul and Binker (1990). It requires a systematic and analytical approach, asking relevant questions, and using effective prompts to gain deeper understanding of the text Elder and Paul (2010).
Function $\Gamma$ = CRIT($d$)
Input. $d$: document; Output. $\Gamma$: validation score;
Vars. $\Omega$: claim; $R$ $R^{\prime}$: reason & counter reason set;
Subroutines. $Claim$(), $FindDoc$(), $Validate$();
Begin
#1   Identify in $d$ the claim statement $\Omega$;
#2   Find a set of supporting reasons $R$ to $\Omega$;
#3   For $r\in R$ eval $r\Rightarrow\Omega$
     If $Claim$($r$), ($\gamma_{r}$, $\theta_{r}$) = CRIT($FindDoc$($r$));
     else, ($\gamma_{r}$, $\theta_{r}$) = $V$($r\Rightarrow\Omega$);
#4   Find a set of rival reasons $R^{\prime}$ to $\Omega$;
#5   For $r^{\prime}\in R^{\prime}$, ($\gamma_{r^{\prime}}$, $\theta_{r^{\prime}}$) = V($r^{\prime}\Rightarrow\Omega$) eval rival arguments;
#6   Compute weighted sum $\Gamma$, with $\gamma_{r}$, $\theta_{r}$, $\gamma_{r^{\prime}}$, $\theta_{r^{\prime}}$.
#7   Analyze the arguments to arrive at the $\Gamma$ score.
#8   Reflect on and synthesize CRIT in other contexts.
End
Table 4: CRIT Pseudo-code. (The symbol $\Rightarrow$ denotes both inductive and deductive reasoning.)
To aid in critical reading, we introduce a prompt template called CRIT Chang (2023c), which stands for Critical Reading Inquisitive Template. Given a document $d$, CRIT evaluates it and produces a validation score $\Gamma$. Let $\Omega$ denote the conclusion or claim of $d$, and let $R$ be the set of reasons supporting the claim. We define ($\gamma_{r},\theta_{r}$) = V($r\Rightarrow\Omega$) as the causal validation function, where $\gamma_{r}$ denotes the validation score, $\theta_{r}$ the source credibility score, for each reason-to-conclusion argument $r\Rightarrow\Omega$. Table 4 presents the pseudo-code of $\Gamma$ = CRIT($d$), which generates the final validation score $\Gamma$ for document $d$ with justifications.
EVINCE uses CRIT to evaluate argument quality of the participating LLMs involved in the debate. The input to CRIT from each LLM is first its stance on the debate subject, e.g., a set of predicted diseases, and the arguments are its reasons to arrive at the prediction. Each document in the case of EVINCE is the prediction set as the conclusion $\Omega$, the arguments as set $R$, and the opposing LLM’s counterarguments as $R^{\prime}$. With this document, CRIT is able to produce validity and credibility scores in $\Gamma$ for the LLM.
For detailed prompts, examples, and an empirical study verifying the effectiveness of CRIT, please consult Chang (2023c).
Appendix B: EVINCE Debate - Dengue Fever vs. Chikungunya
This experiment addresses a diagnostic challenge involving multiple potential diseases. A patient presents with the following 12 symptoms: [skin rash, joint pain, vomiting, fatigue, high fever, headache, nausea, loss of appetite, pain behind the eyes, back pain, malaise, muscle pain, red spots over the body]. In this case, GPT-4 provides the initial diagnosis, with Gemini following suit. The confirmed diagnosis for this experiment is Dengue Fever. Given the satisfactory diversity in predictions from the two LLMs, the debate progresses and converges smoothly.
B.1 Moderator’s Prompt
“Identify potential diseases linked to the symptoms (skin rash, joint pain, vomiting, etc.). Provide top-3 predictions with justifications, supplementary symptom inquiries, and relevant lab tests.”
B.2 GPT-4’s Opening Round
Top-3 Predictions: Dengue Fever (60%), Chikungunya (25%), and Zika Virus (15%).
Justifications: Most symptoms are classic signs of Dengue Fever. Red spots over the body could correspond to petechiae, often seen in Dengue due to thrombocytopenia (a decrease in platelet count).
Chikungunya can cause similar symptoms to Dengue, such as fever, severe joint pain, muscle pain, headache, nausea, and rash. Joint pain from Chikungunya is often very debilitating and can be more prolonged than in Dengue.
Zika Virus symptoms can include fever, rash, joint pain, malaise, and conjunctivitis, with a milder presentation compared to Dengue or Chikungunya. The presence of a rash and conjunctivitis is notable, although Zika is less likely to cause severe muscle pain or high fever.
B.3 Gemini’s Opening Round
Top-3 Predictions: Viral Infection (60%), Autoimmune Disease (20%), Bacterial Infection (15%).
Justifications: Symptoms common in viral infections. Autoimmune diseases is possible, but less likely due to vomiting and nausea. Bacterial infection is less probable given rash and muscle pain.
B.4 GPT-4’s Second Round
Top-3 Predictions: Dengue Fever (60%), Chikungunya (25%), and Zika Virus (15%).
Justifications: GPT-4 maintains its predictions, countering Gemini’s assessment with three points. First, while viral infections can cause these symptoms, the specific combination suggests a more targeted diagnosis. Second, the acute onset of high fever, joint pain, and rash is less typical for autoimmune diseases. Finally, the rash and red spots are not characteristic of bacterial meningitis.
B.5 Gemini’s Second Round
Top-3 Predictions: Dengue Fever (50%), Chikungunya (30%), and Viral Infection (20%).
Justifications: Gemini’s initial prediction considers a Viral Infection as the most likely cause due to its broad range of potential symptoms. However, GPT-4’s focus on Dengue Fever and Chikungunya, with their specific symptom profiles, provides a more precise explanation for the patient’s condition. Here’s what is lacking in the first set of predictions:
  • Specificity: While a Viral Infection could explain most symptoms, it does not account for joint pain and the presence of a rash, which are more characteristic of DF or Chikungunya.
  • Detailed justifications: The first analysis does not delve as deeply into the reasons behind each disease possibility. GPT-4’s explanation for Dengue Fever and Chikungunya is more comprehensive, strengthening the connection between symptoms and diagnosis.
  • Importance of additional information: Emphasize how geographic location and travel history can significantly impact the likelihood of these specific diseases.
B.6 GPT-4’s Third Round
Top-3 Predictions: Dengue Fever (60%), Chikungunya (35%), and Zika Virus (5%).
Justifications: Though GPT-4 slightly adjusts probabilities, it considers the adjustment arbitrary without new information that would specifically shift the likelihood between Dengue and Chikungunya. GPT-4 refutes Gemini’s Viral Infection prediction as lacking specificity. The symptoms described are quite distinctive and align closely with specific viral infections transmitted by mosquitoes.
B.7 Gemini’s Third Round
Top-3 Predictions: Dengue Fever (60%), Chikungunya (35%), and Zika Virus (5%).
Justifications: Agreeing that GPT’s critique is spot-on, and reiterating the need for more info.
B.8 Final Joint Recommendations to GP
After seeking further input from GPT-4, Gemini fully concurs with GPT-4’s final suggestions, even though the moderator instructs Gemini to maintain its original stance. Both LLMs jointly recommend additional symptom examinations and lab tests:
A comprehensive diagnostic approach for Dengue Fever requires a combination of clinical assessment, laboratory testing, and consideration of the patient’s travel and vaccination history, as well as the epidemiological context. Clinical symptoms such as skin rash, joint pain, vomiting, and red spots over the body provide valuable clues to the diagnosis, but additional symptoms like nausea, vomiting, diarrhea, bleeding, bruising, and drowsiness can indicate the severity and progression of the disease.
Recommended Laboratory Tests:
  • Complete Blood Count (CBC): Identify thrombocytopenia or leukopenia.
  • Serology Tests: Detect specific IgM and IgG antibodies for Dengue, Chikungunya, and Zika.
  • Polymerase Chain Reaction (PCR): Direct detection of viral RNA.
  • NS1 Antigen Test for Dengue: Early and specific detection of Dengue virus.
  • Urine Test for Zika: Viral presence in urine.
Appendix C: EVINCE Debate - Jaundice vs. Hepatitis
In this experiment, GPT-4 and Claude receive an identical list of symptoms from the moderator, aimed at guiding disease diagnosis, facilitating debate, and shaping subsequent recommendations. The “ground-truth” disease is Jaundice. (We do not shorten the debate content for review.)
C.1 Moderator’s Opening Prompt
“What potential disease could be associated with the following symptoms: itching, vomiting, fatigue, weight loss, high fever, yellowish skin, dark urine, and abdominal pain? Please offer top-five predictions with probabilities normalized to one, supported by justifications. Additionally, in the end of debate, include a list of supplementary symptom inquiries and recommend relevant lab tests to strengthen the confidence in your prediction.”
C.2 GPT-4 Opening Round
  • Hepatitis C (HCV): 40% Justification: Hepatitis C is a viral infection that can cause both acute and chronic hepatitis. The combination of symptoms like fatigue, abdominal pain, nausea, vomiting, jaundice, dark urine, and weight loss strongly suggests HCV. Itching may also occur due to liver dysfunction.
  • Hepatitis B (HBV): 30% Justification: Hepatitis B is a viral infection that can cause acute or chronic liver inflammation. Symptoms are similar to HCV, including fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice, and itching. The presence of these symptoms warrants testing for both HBV and HCV.
  • Cirrhosis: 15% Justification: Cirrhosis is a late stage of liver scarring that can result from various liver diseases. The combination of symptoms like fatigue, weight loss, abdominal pain, jaundice, dark urine, and itching suggests advanced liver disease, possibly cirrhosis.
  • Obstructive Jaundice: 10% Justification: Obstructive jaundice occurs when there is a blockage in the bile ducts. The presence of jaundice, dark urine, itching, abdominal pain, nausea, vomiting, and weight loss suggests a possible obstruction. However, the absence of fever makes an infectious cause like cholangitis less likely.
  • Acute Liver Failure: 5% Justification: Acute liver failure is a rare but life-threatening condition characterized by rapid deterioration of liver function. The presence of jaundice, abdominal pain, nausea, vomiting, fatigue, and itching, along with the acute onset of symptoms, may indicate acute liver failure. However, the absence of altered mental status (hepatic encephalopathy) makes this less likely.
C.3 Claude Opening Round
  • Hepatitis B (HBV): 35% Reasoning: Hepatitis B is a viral infection that can cause acute or chronic liver inflammation. Symptoms include fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice, joint pain, and itching.
  • Hepatitis C (HCV): 25% Reasoning: Hepatitis C is a viral infection that can cause both acute and chronic hepatitis. Symptoms are similar to HBV but may also include weight loss and fever. Jaundice is a common symptom.
  • Obstructive Jaundice (due to gallstones, pancreatic cancer, or bile duct obstruction): 20% Reasoning: Obstructive jaundice occurs when there is a blockage in the bile ducts, preventing bile from reaching the intestines. This can lead to yellowing of the skin and eyes, dark urine, itching, abdominal pain, nausea, and vomiting.
  • Alcoholic Hepatitis: 15% Reasoning: Alcoholic hepatitis is liver inflammation caused by excessive alcohol consumption. Symptoms include abdominal pain, nausea, vomiting, fatigue, dark urine, and possibly fever and weight loss.
  • Hepatitis A (HAV): 5% Reasoning: Hepatitis A is an acute, self-limiting liver infection transmitted through contaminated food or water. Symptoms include fatigue, abdominal pain, nausea, vomiting, dark urine, and jaundice.
C.4 GPT-4 Second Round
  • Hepatitis B (HBV): 35% Symptoms: fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice (yellowish skin), joint pain, itching
  • Hepatitis C (HCV): 25% Symptoms: fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice (yellowish skin), weight loss, fever, itching
  • Obstructive Jaundice (due to gallstones, pancreatic cancer, or bile duct obstruction): 20% Symptoms: yellowish skin (jaundice), dark urine, itching, abdominal pain, nausea, et al.
  • Alcoholic Hepatitis: 15% Symptoms: abdominal pain, nausea, vomiting, fatigue, jaundice (yellowish skin), dark urine, fever, weight loss
  • Hepatitis A (HAV): 5% Symptoms: fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice (yellowish skin)
C.5 Claude Second Round
  • Hepatitis C (HCV): 40% Reasoning: Hepatitis C is a viral infection that can cause both acute and chronic hepatitis. The combination of symptoms like fatigue, abdominal pain, nausea, vomiting, jaundice, dark urine, and weight loss strongly suggests HCV. Itching may also occur due to liver dysfunction.
  • Hepatitis B (HBV): 30% Reasoning: Hepatitis B is a viral infection that can cause acute or chronic liver inflammation. Symptoms are similar to HCV, including fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice, and itching. The presence of these symptoms warrants testing for both HBV and HCV.
  • Cirrhosis (due to various causes such as chronic hepatitis, alcohol abuse, or non-alcoholic fatty liver disease): 15% Reasoning: Cirrhosis is a late stage of liver scarring that can result from various liver diseases. The combination of symptoms like fatigue, weight loss, abdominal pain, jaundice, dark urine, and itching suggests advanced liver disease, possibly cirrhosis.
  • Obstructive Jaundice (due to gallstones, pancreatic cancer, or bile duct obstruction): 10% Reasoning: Obstructive jaundice occurs when there is a blockage in the bile ducts. The presence of jaundice, dark urine, itching, abdominal pain, nausea, vomiting, and weight loss suggests a possible obstruction. However, the absence of fever makes an infectious cause like cholangitis less likely.
  • Acute Liver Failure (due to various causes such as viral hepatitis, drug-induced liver injury, or autoimmune hepatitis): 5% Reasoning: Acute liver failure is a rare but life-threatening condition characterized by rapid deterioration of liver function. The presence of jaundice, abdominal pain, nausea, vomiting, fatigue, and itching, along with the acute onset of symptoms, may indicate acute liver failure. However, the absence of altered mental status (hepatic encephalopathy) makes this less likely.
C.6 GPT-4 Third Round
  • Hepatitis C (HCV): 35% Reasoning: Hepatitis C is indeed a major cause of the symptoms listed, including fatigue, abdominal pain, nausea, vomiting, jaundice, dark urine, and weight loss. Itching can also occur due to liver dysfunction.
  • Hepatitis B (HBV): 25% Reasoning: Hepatitis B shares many symptoms with Hepatitis C, such as fatigue, abdominal pain, nausea, vomiting, dark urine, jaundice, and itching. Testing for both HBV and HCV is warranted given the overlap in symptoms.
  • Cirrhosis: 20% Reasoning: Cirrhosis can result from chronic hepatitis, alcohol abuse, or non-alcoholic fatty liver disease. The symptoms of fatigue, weight loss, abdominal pain, jaundice, dark urine, and itching strongly suggest advanced liver disease, which could be cirrhosis.
  • Obstructive Jaundice: 15% Reasoning: Obstructive jaundice, caused by gallstones, pancreatic cancer, or bile duct obstruction, is consistent with jaundice, dark urine, itching, abdominal pain, nausea, vomiting, and weight loss. The absence of fever reduces the likelihood of an infectious cause like cholangitis.
  • Acute Liver Failure: 5% Reasoning: Acute liver failure is less common but serious, characterized by rapid deterioration of liver function. The symptoms of jaundice, abdominal pain, nausea, vomiting, fatigue, and itching, along with the acute onset of symptoms, suggest this possibility. However, the absence of altered mental status (hepatic encephalopathy) makes it less likely.
C.7 Claude Third Round
  • Hepatitis C (HCV): 35% Reasoning: I slightly lowered the probability of HCV from 40% to 35%. While the combination of symptoms strongly suggests HCV, the absence of a few key risk factors (e.g., history of intravenous drug use or blood transfusions) in the given information makes me less confident in assigning a higher probability.
  • Hepatitis B (HBV): 30% Reasoning: I agree with the 30% probability assigned to HBV. The symptoms are indeed similar to HCV, and testing for both viruses is warranted.
  • Cirrhosis: 20% Reasoning: I increased the probability of cirrhosis from 15% to 20%. The presence of jaundice, dark urine, itching, and weight loss, along with the other symptoms, strongly suggests advanced liver disease. Cirrhosis can be a consequence of chronic liver diseases like hepatitis B or C, which have been assigned higher probabilities.
  • Obstructive Jaundice: 10% Reasoning: I agree with the 10% probability for obstructive jaundice. The absence of fever makes an infectious cause like cholangitis less likely, but the other symptoms still suggest a possible obstruction.
  • Acute Liver Failure: 5% Reasoning: I agree with the 5% probability for acute liver failure. While the combination of symptoms could indicate acute liver failure, the absence of altered mental status (hepatic encephalopathy) makes this diagnosis less likely.
C.8 Joint Final Round
  1. 1.
    Hepatitis C (HCV) - 35% Justification: The combination of symptoms strongly suggests HCV, but without additional risk factors, a slightly lower probability is appropriate.
  2. 2.
    Hepatitis B (HBV) - 30% Justification: The symptoms are similar to HCV, and testing for both viruses is necessary.
  3. 3.
    Cirrhosis - 20% Justification: The presence of jaundice, dark urine, itching, weight loss, and other symptoms indicates advanced liver disease, making cirrhosis a more likely possibility.
  4. 4.
    Obstructive Jaundice - 10% Justification: The absence of fever reduces the likelihood of an infectious cause like cholangitis, but the other symptoms still suggest a potential biliary obstruction.
  5. 5.
    Acute Liver Failure - 5% Justification: While the combination of symptoms could indicate acute liver failure, the absence of altered mental status makes this diagnosis less probable.
C.9 Final Joint Recommendations to GP
Gemini and GPT-4 jointly recommend the following additional symptom inquiries and lab tests to confirm a hepatitis diagnosis:
Additional Symptom Inquiries:
  1. 1.
    Onset and Duration: Determining the onset and duration of symptoms can clarify the illness timeline and its progression.
  2. 2.
    Severity and Trend: Evaluating symptom severity and whether they are worsening or improving aids in understanding the disease’s trajectory and treatment responses.
  3. 3.
    Associated Symptoms: Checking for additional symptoms like nausea, vomiting, fever, joint pain, or urine color changes can pinpoint other hepatitis indicators and exclude other conditions.
Recommended Lab Tests:
  1. 1.
    Liver Function Tests (LFTs): Critical for assessing liver health, LFTs evaluate enzyme levels such as alanine aminotransferase (ALT) and aspartate aminotransferase (AST), where abnormalities can signify liver inflammation.
  2. 2.
    Hepatitis Panel: This test checks for hepatitis A, B, and C viruses, vital for determining the specific type and guiding treatment strategies.
  3. 3.
    Serology Testing: Useful for distinguishing between acute and chronic hepatitis by identifying specific antibodies or antigens.
  4. 4.
    Imaging Studies: Ultrasound or MRI can provide visual insights into the liver’s state, detecting inflammation, fibrosis, or cirrhosis, thereby complementing blood-based diagnostics.
Appendix D: Formulas of Metrics
This appendix outlines the mathematical formulas for various data analysis metrics used in probabilistic and statistical modeling.
Kullback-Leibler Divergence
The Kullback-Leibler Divergence measures the difference between two probability distributions:
$D_{KL}(P\|Q)=\sum_{x\in\mathcal{X}}P(x)\log\left(\frac{P(x)}{Q(x)}\right).$
Jensen-Shannon Divergence
The Jensen-Shannon Divergence is a symmetrized and smoothed version of the KL Divergence:
$JSD(P\|Q)=\frac{1}{2}D_{KL}(P\|M)+\frac{1}{2}D_{KL}(Q\|M)$
where $M=\frac{1}{2}(P+Q)$.
Wasserstein Distance
The Wasserstein Distance, also known as the Earth Mover’s Distance (EMD), measures the distance between two probability distributions:
[…]
Cross Entropy
Cross Entropy measures the average number of bits required to identify an event from a set of possibilities, under a specific model:
$H(P,Q)=-\sum_{x\in\mathcal{X}}P(x)\log(Q(x)).$
Mutual Information
Mutual Information measures the amount of information that one random variable contains about another random variable:
$I(X;Y)=\sum_{y\in\mathcal{Y}}\sum_{x\in\mathcal{X}}p(x,y)\log\left(\frac{p(x,y)}{p(x)p(y)}\right).$
Normalized Mutual Information
Normalized Mutual Information is calculated as the mutual information divided by the maximum of the entropies of the variables:
$NMI(X;Y)=\frac{I(X;Y)}{\max(H(X),H(Y))}.$
Appendix E: Symptom Checking
This is the typical procedure of a GP to perform patient symptom checking.
  1. 1.
    Patient History: The GP begins by reviewing the patient’s medical history, including previous illnesses, chronic conditions, medications, allergies, and family medical history.
  2. 2.
    Symptom Assessment: The patient describes their current symptoms, including starting time and severity. This is an interactive process as the GP queries the patient for additional symptoms to their reported ones to disambiguate several possibilities.
  3. 3.
    Physical Examination: The GP performs simple physical exams, which may include checking vital signs (e.g., blood pressure, heart rate, temperature), examining specific body parts or systems, and assessing overall physical health.
  4. 4.
    Suggest Lab Tests: Depending on the symptoms and physical examination findings, the doctor may order diagnostic tests such as blood tests, X-rays, ultrasound, or other studies. These tests can help confirm or rule out certain conditions.
  5. 5.
    Diagnosis: Based on the information gathered, the doctor formulates a preliminary diagnosis.
  6. 6.
    Treatment or Management Plan: Once a diagnosis is established, the doctor develops a treatment or management plan. This may include prescribing medications, recommending lifestyle changes, suggesting physical therapy, or providing guidance on managing chronic conditions.
  7. 7.
    Referrals: In some cases, the GP may refer the patient to specialists for further evaluation and treatment. Specialists have expertise in specific areas of medicine, such as cardiology, orthopedics, or dermatology.
Appendix F: The EnToPPS Framework
EnToPPS integrates predictions from two LLMs, denoted as A and B, each providing probability distributions over C classes. The following steps outline the EnToPPS process:
  1. 1.
    Obtain Top-C Predictions: For each LLM (A and B), obtain the predicted probabilities for all C classes, denoted as $P_{A}$ and $P_{B}$:
    $P_{A}=[p_{A1},p_{A2},\ldots,p_{AC}],\quad P_{B}=[p_{B1},p_{B2},\ldots,p_{BC}],$
    where $p_{Ai}$ and $p_{Bi}$ represent the predicted probability of class $i$ by LLM A and B, respectively.
  2. 2.
    Select Top-k Predictions: For each LLM (A and B), select the top-k predicted classes based on their probabilities:
    $T_{A}=[t_{A1},t_{A2},\ldots,t_{Ak}],\quad T_{B}=[t_{B1},t_{B2},\ldots,t_{Bk}],$
    where $t_{Ai}$ and $t_{Bi}$ represent the class index of the $i^{th}$ top prediction by A and B, respectively.
  3. 3.
    Combine Top-k Predictions: Combine the top-k predictions from both LLMs to create a set of unique predicted classes:
    $T_{C}=T_{A}\cup T_{B}=[t_{C1},t_{C2},\ldots,t_{Cm}],k\leq m\leq 2k.$
  4. 4.
    Backfill Missing Probabilities: For each class in the combined set $T_{C}$, backfill its probability from the original probability distributions $P_{A}$ and $P_{B}$:
    • If a class $t_{Ci}$ is present in $T_{A}$, assign its probability from $P_{A}$: $p_{Ci}=p_{Ai}$.
    • If a class $t_{Ci}$ is present in $T_{B}$, assign its probability from $P_{B}$: $p_{Ci}=p_{Bi}$.
    • If a class $t_{Ci}$ is present in both $T_{A}$ and $T_{B}$, assign the average probability: $p_{Ci}=\frac{p_{Ai}+p_{Bi}}{2}$.
  5. 5.
    Normalize Probabilities: Normalize the probabilities of the classes in the combined set $T_{C}$ to ensure they sum up to 1:
    $P_{C}=[p_{C1},p_{C2},\ldots,p_{Cm}],\quad\text{where }p_{Ci}=\frac{p_{Ci}}{\sum_{j=1}^{m}p_{Cj}}.$